Calculate the volume occupied by 0.8 moles of a gas at 27 degree celsius temperature and atmospheric pressure if the compressibility factor of a gas is 0.86

Grade:12th pass

1 Answers

Susmita
5 years ago
Compressibility factor z=pV/nRT
Here z=0.86,n=0.8,p=1atm,T=27+273=300 K,R=0.082lt atm/mole/K.
So V=znRT/p=16.92 lt.
If it helps please approve.